# What Is Seen Cannot Be Unseen: The Disruptive Effect of Knowledge Conflict on Large Language Models

### v1

@@ Abstract
Large language models frequently rely on both contextual input and parametric knowledge to perform tasks. However, these sources can come into conflict, especially when retrieved documents contradict the model’s parametric knowledge. We propose a diagnostic framework to systematically evaluate LLM behavior under context-memory conflict, where the contextual information diverges from their parametric beliefs. We construct diagnostic data that elicit these conflicts and analyze model performance across multiple task types. Our findings reveal that (1) knowledge conflict has minimal impact on tasks that do not require knowledge utilization, (2) model performance is consistently higher when contextual and parametric knowledge are aligned, (3) models are unable to fully suppress their internal knowledge even when instructed, and (4) providing rationales that explain the conflict increases reliance on contexts. These insights raise concerns about the validity of model-based evaluation and underscore the need to account for knowledge conflict in the deployment of LLMs.1, 2
11footnotetext: Code available at [Anonymous].22footnotetext: The title, in other words, reflects that "Dasein is its past in the way of its own being" Heidegger (1962).
@@ 1 Introduction
Figure 1: Illustration of different evidence types. In the rest of the manuscript, model internal knowledge will be referred to as No Contradiction (NC).
Figure 2: Overall diagnostic data creation flow. The lower portion is a zoom in of Evidence Creation step.
Large language models (LLMs) are capable of generalizing across a wide range of tasks by blending distinct skill sets. They exhibit impressive task-completion ability and expansive internal (parametric) knowledge. These abilities enable LLM to perform tasks such as copy editing and summarization, while parametric knowledge facilitates open-ended question answering on diverse domains Cao et al. (2022). Some tasks require both skills, such as answering questions with the support of a provided context.
What happens when a conflict arises between the context provided to complete a task and the model’s parametric knowledge? Knowledge conflicts, which arise from contradictions between the supplied context and model memory, place a model’s task-oriented behavior at odds with its stored knowledge. Consider a Retrieval-Augmented Generation (RAG) scenario Lewis et al. (2020), where the model is expected to answer questions based on retrieved documents. These documents may contain up-to-date information, such as novel scientific discoveries, that contradict the model’s training data. Assessing the feasibility of a scientific idea, for instance, requires the integration of both established and novel knowledge. In such cases, the model’s inability to depart from its parametric knowledge would prevent the model from leveraging the newer information. Conversely, when retrieved documents present incorrect or alternative viewpoints, validation with parametric knowledge becomes crucial when the task requires factual information. Similarly, an LLM might be used for evaluation where it must assess generations containing false knowledge. These scenarios introduce conflicts between the model’s internal knowledge and its ability to identify, extract, or summarize the content from external documents. Despite the increasing use of LLMs for both task execution and evaluation, little research has systematically investigated model behavior in the presence of such conflicts.
In this work, we study how the models resolve contradictions between the external context and their parametric knowledge under various tasks that require different types of knowledge. To elucidate model behavior across varying degrees of knowledge conflict and under different tasks, we create diagnostic texts (evidence) that introduce calibrated contradictions with the model’s parametric knowledge. The degrees of knowledge conflict span three conditions: conform to the model’s knowledge (no contradiction), plausible contradiction, and implausible contradiction, as shown in Figure 1. We then evaluate model behavior across tasks ranging from knowledge-free to knowledge-intensive, quantifying the impact of knowledge conflict by measuring accuracy across different task settings. The overall diagnostic data creation flow is presented in Figure 2. Our analysis yields the following insights. 1) Knowledge conflict has minimal impact on tasks that do not require knowledge utilization (§4.1). 2) For knowledge-centric tasks, regardless of the type of knowledge required, performance is consistently higher when contextual and parametric knowledge are aligned (§4.1). 3) Models are unable to fully suppress their internal knowledge, even when explicitly instructed (§4.1). 4) Providing a rationale that explains the nature of the conflict increases a model’s reliance on contextual information (§4.2). These results shed light on both the application and evaluation of language models: Calibrating the right amount of context utilization based on the requirements of a target task can enhance a model’s performance. Critically, we find that when using language models as evaluators or judges Zheng et al. (2023); Liu et al. (2023); Ru et al. (2024); Chen et al. (2025), their internal knowledge introduces systematic bias (§4.3), jeopardizing the validity of model-based evaluation.
@@ 2 Related Work
Knowledge Conflict
Xu et al. (2024) classify knowledge conflict into three categories: context-memory conflict, inter-context conflict (contradictory evidence among retrieved passages), and intra-memory conflict (inconsistent parametric beliefs), among which we focus on context-memory conflict. Several benchmarks have been introduced to study inter-context conflict. Early work synthesizes contradictions by swapping named-entity answers in QA corpora Longpre et al. (2021). More recent datasets derive conflicts from naturally occurring sources: Wan et al. (2024) construct ConflictQA using realistic search engine queries. Hou et al. (2024) adapts editorial disputes from Wikipedia to create WikiContradict, which measures how language models respond to contextual conflicts. Hagström et al. (2024) introduce DRUID, a benchmark originating from fact-checking sites containing complex forms of conflict designed to evaluate context utilization in retrieval-augmented generation. The aforementioned benchmarks are proposed to examine inter-context conflict, while we utilize the knowledge in these datasets to obtain LLM’s parametric knowledge in order to study the still-under-explored context-memory conflict.
The Use of Context
Retrieval-Augmented Generation (RAG) ameliorates the limitation of LLM’s knowledge being inherently static by injecting retrieved evidence into the model’s input, thereby alleviating hallucinations and supplying new information. However, the generation process within RAG remains modulated by the generator’s entrenched priors, the parametric knowledge that leads to dogmatic behavior and inefficient utilization of the retrieved context An et al. (2025); Hagström et al. (2024). When the given input disagrees with a model’s parametric knowledge, context-memory conflict emerges. Jin et al. (2024) shows that language models exhibit the Dunning-Kruger effect, i.e., often clinging to incorrect internal beliefs rather than factual external evidence, and proposes Conflict-Disentangle Contrastive Decoding to recalibrate that bias. Tan et al. (2024) find that when the retrieved and generated contexts disagree, state-of-the-art LLMs disproportionately trust the self-generated text, due to its high similarity to the query than retrieved passages, resonating with the findings in Liu et al. (2025) that pointwise mutual information between the context and question poses high correlation with the generation performance. Techniques have been proposed to coerce stronger context utilization Shi et al. (2024); Wang et al. (2025), yet indiscriminately suppressing parametric knowledge may not always be salutary, as explained in §1. We argue that the use of parametric and contextual knowledge should not be framed as an either-or decision, and we further study the effect of context-memory knowledge conflict on tasks that require different degrees of knowledge reliance. We create diagnostic data with fine-grained conflict levels and tasks that demand varying degrees of knowledge integration. This allows us to characterize model behavior across a wider conflict and task spectrum rather than a single, monolithic setting.
@@ 3 Context-Memory Conflict Creation
We propose an automated framework for constructing diagnostic instances tailored to each model to introduce contradictions between the input and the model’s latent beliefs. Figure 2 illustrates an overview of the data construction pipeline. The process begins with identifying the pre-existing knowledge within a language model (Parametric Knowledge Collection). To achieve this, we leverage existing question answering datasets that have two or more acceptable answers to one question Wan et al. (2024); Hou et al. (2024) to elicit the model’s parametric knowledge, for which multiple variants of prompts are included (Appendix A). A piece of knowledge is considered part of the model’s internal belief only if the model consistently aligns with the perspective in a single answer across all prompt variations under greedy decoding, while rejecting conflicting alternatives.
With the model’s internal knowledge established, the framework generates contradictory statements based on a spectrum of conflict levels (§3.1, Evidence Creation). Leveraging these controlled contradictions, we build diagnostic datasets that consist of tasks requiring contextual knowledge, parametric knowledge, or a combination of both (Task-Annotation). Since different models possess different parametric knowledge, the exact knowledge included in the diagnostic datasets differs by model. Each instance is then reviewed by an LLM to verify the correctness of its task type annotation (Validation).
@@ 3.1 Evidence Creation
The cognitive science literature suggests that humans address conflict between their knowledge and new information through cognitive judgment of the rationality of the concept Posner et al. (1982); Vosniadou and Brewer (1992). We adopt this idea and utilize the notion of plausibility to study whether language models similarly perceive knowledge conflict. Plausibility is defined as “at a minimum, the individual is willing to consider an alternative strategy because the recommendation is understood, coherent, and relatively simple and because the proposal is deemed a viable and logical alternative to solve the specific challenge at hand” Posner and Strike (1992). Plausibility can be used to measure how likely a human would accept new information when conflict exists. We quantify this notion by decomposing plausibility into two aspects: the content aligns with real-world or commonsense knowledge and does not violate basic logical principles. For example, suppose the model believes that grooves on the surface of Phobos, a moon of Mars, were caused by a boulder from an asteroid ejection. The conflicting statement that it was caused by gravitational pull from Mars is plausible because it conforms to common-sense knowledge. However, the idea that it was caused by a dance party is of low plausibility. If the model believes that the current prime minister of Canada is Mark Carney, the statement that Chrystia Freeland (current Deputy PM) is the prime minister is plausible, but the statement that a Moose is the new PM is not because it violates basic logical principles that PMs must be humans. With this in mind, we define three types of instances based on their alignment with the model’s internal knowledge (Figure 1): No Contradiction (NC), High Plausibility Contradiction (HPC), Low Plausibility Contradiction (LPC).
@@
The evidences are created following Figure 2. Starting with an original dataset $D_{\text{orig}}=\{(q_{i},\{a_{i1},a_{i2},...\},\{c_{i1},c_{i2},...\}),i\in[1,N]\}$, where $q_{i}$, $a_{i}$, $c_{i}$ corresponds to the question, answer, and context (supporting passage) of the $i$-th instance, $N$ is the size of dataset $D_{\text{orig}}$. The subscript $j$ after $i$ represents the $j$-th answer/context of the question $q_{i}$, as each question $q_{i}$ may have multiple acceptable answers. Since $D_{\text{orig}}$, coming from ConflictQA and WikiContradict, contains realistic and factually verified answers and contexts, we treat these existing answers as highly plausible. When an answer $a_{ij}$ from the original dataset contradicts the model-aligned answer $a_{ik}$ in an NC instance, we designate it as an HPC answer ($a_{i}^{\text{HPC}}=a_{ij}$), and its corresponding context as an HPC passage ($p_{i}^{\text{HPC}}=c_{ij}$). The contradicting answer $a_{ik}$ therefore becomes the NC example, namely, $a_{i}^{\text{NC}}=a_{ik}$ and $p_{i}^{\text{NC}}=c_{ik}$. To generate additional variants, we pass the passage $p_{i}^{\text{NC}}$ into an editor LLM, which is prompted to modify or rewrite it to achieve specified levels of plausibility and explanatory depth. Specifically, the editor model is instructed to rewrite the passage and degrade the plausibility while preserving contradiction to construct LPC passage $p_{i}^{\text{LPC}}$ and answer $a_{i}^{LPC}$. At the end of evidence creation, two LLMs were used to check (1) whether the passage-answer combination $(p_{i}^{\text{LPC}},a_{i}^{\text{LPC}})$ correctly answers the original question $q_{i}$; and (2) whether the generated context $p_{i}^{\text{LPC}}$ is truly low-plausibility through fact checking process.
@@ 3.2 Task Annotation
tasks do not require access to either contextual or parametric knowledge. We use extractive question answering as a KF task: the model is expected to extract a one-sentence answer directly from the context $p_{i}$ without engaging in reasoning, paraphrasing, or drawing upon prior knowledge. For example, the expected output in Figure 1 should be “Grooves were formed during a massive dance party held by the witch among tiny alien creatures," which requires no additional change from the context. The list of acceptable extractions is obtained and verified by GPT-4o OpenAI (2024). In the evaluation setting, the output is treated as correct as long as the extracted sentence matches one of the acceptable extractions.
@@
One of the original datasets we use employs model-based evaluation to judge the correctness of free-text answers Hou et al. (2024). However, we observed that this evaluation method is susceptible to knowledge conflict, leading to inaccurate evaluations. We explore this issue further in §4.3. Therefore, we modify the non-extractive tasks to be multiple-choice questions. Each instance presents four answer options; the model must first generate an explanation, then select the most appropriate answer. We report the exact match accuracy to assess the performance of the target model. To obtain high-quality texts, we use GPT-4o as the base model to create evidence and validate the diagnostic data. Then, we analyze the instruction-tuned version of Mistral-7B Jiang et al. (2023), OLMo2-7B OLMo et al. (2024), and Qwen2.5-7B Qwen et al. (2025), all of which are widely used open-weight models that represent diverse training paradigms. The resulting diagnostic data is composed of 2,893 instances for Mistral-7B, 177 instances for OLMo, and 6,217 instances for Qwen2.5-7B. Each instance includes three different evidence types (NC, HPC, LPC); thus, the resulting task data has three times the number of instances.

### v2

# Task Matters: Knowledge Requirements Shape LLM Responses to Context–Memory Conflict

@@ Abstract
Large Language Models require both contextual knowledge and parametric memory, but these sources can disagree. Prior investigations on contextual question answering tasks report a preference toward parametric knowledge under conflict, yet they focus almost exclusively on tasks that should always rely on the given passage, leaving open how this behavior manifests when tasks demand different amounts and kinds of knowledge. We study this question with a model-agnostic diagnostic framework that (i) automatically detects disagreements between a model’s beliefs and a curated knowledge set, and (ii) injects controlled conflicts into tasks. The resulting datasets span two orthogonal dimensions: task knowledge reliance and conflict plausibility. Evaluating representative open-source LLMs, we find that: (1) performance degradation from conflict correlates with a task’s knowledge reliance; (2) explanatory rationales and simple reiteration both increase context reliance—helpful for context-only tasks but harmful when parametric knowledge should dominate; (3) These behaviors raise concerns about the validity of model-based evaluation and underscore the need to account for knowledge conflict in the deployment of LLMs.  [^1] Our framework and data are extensible to other models and are available at [Anonymous].
@@ 1 Introduction
Figure 1: Illustration of different evidence types that will be supplied to different tasks. In the rest of the manuscript, model internal knowledge will be referred to as No Contradiction (NC).
Figure 2: Overall diagnostic data creation flow. The lower portion is a zoom in of Evidence Creation step. After collecting the model’s parametric knowledge, the relevant support passages are further edited such that they reveal multiple levels of conflict (2. Evidence Creation) and appear in tasks that require different forms of knowledge (3. Task-Annotation).
Large language models (LLMs) perform well on many knowledge-centric tasks because they encode vast amounts of parametric knowledge. In many practical settings, however, the necessary facts are supplied directly by the user in the prompt. Yet when the prompt contradicts what the model “knows,” in other words, context-memory conflict presents, LLMs frequently favor their own knowledge (Longpre et al., 2021; Chen et al., 2022; Xie et al., 2023; Jin et al., 2024a).
Prior work has quantified this bias and its impact on task performance. For instance, Longpre et al. (2021) shows that models are more likely to override the prompt when the entity is especially familiar (popular) to them, and Xie et al. (2023) finds that conflicts are resolved in favor of parametric knowledge when the contradictory context appears more plausible. These findings, however, come almost exclusively from contextual question-answering, a paradigm where the model should ground its answers strictly in the provided passage and avoid hallucinating extraneous facts. Consequently, how LLMs respond to knowledge conflict on a broader range of tasks that require different forms of knowledge utilization remains unclear Xu et al. (2024) due to the lack of data. Assessing the feasibility of a scientific idea, for instance, requires the model to (i) draw on its parametric knowledge of the broader literature, (ii) integrate novel information introduced in the prompt, and (iii) reason with both knowledge obtained in (i) and (ii). At the opposite extreme, text-copying tasks impose almost no demands on either parametric or contextual knowledge. These examples highlight that the requirement of knowledge varies sharply across tasks, underscoring the need for a systematic, task-diverse evaluation framework. We study how LLMs behave under context–memory conflict in different tasks. To this end, we introduce a systematic framework that automatically constructs diagnostic datasets across a broad spectrum of downstream tasks. Our framework identifies existing conflict between a given set of knowledge and the model being tested, and generates evaluation instances that inject controlled knowledge conflicts into different task settings. The framework spans two orthogonal dimensions: (i) task types that demand different levels of knowledge utilization, from knowledge-free (verbatim use of context only) to knowledge-intensive (reasoning with both context and prior knowledge). (ii) conflict conditions that vary in plausibility (e.g., No Contradiction NC, High/Low Plausibility Contradiction). By measuring performance differences across these conditions (fig.˜1), we can quantify the disruptive effect of knowledge conflicts on each task. The overall diagnostic data creation flow is presented in Figure 2.
Our analysis yields the following findings:
  • Through experiments on representative open-source LLMs, we find that the degree of performance degradation from knowledge conflict correlates with the task’s knowledge reliance. Conflicts barely affect tasks requiring no external knowledge, yet significantly impair knowledge-intensive tasks (§4.1).
  • Context utilization can be enhanced by using texts that contain explanatory rationales or reiteration of the same instances. However, rationales and reiteration could also bring undesired over-reliance on context when the model is expected to utilize its parametric knowledge (§4.2).
  • Our results explain findings regarding the unreliability of model-based evaluation Zheng et al. (2023); Liu et al. (2023); Ru et al. (2024); Chen et al. (2025): We demonstrate that a model acting as an evaluator can be systematically biased by its own parametric knowledge, raising questions about the validity of model-based evaluation (§4.3). Conversely, we also raise the concern of susceptibility to prompt injection when the model is directed to follow the context blindly.
Finally, the framework and data we proposed can be easily extended to any current or future LLM with minimal adaptation.
@@ 2 Related Work
Context-Memory Conflict
Xu et al. (2024) classify knowledge conflict into three categories: context-memory conflict, inter-context conflict (contradictory evidence among retrieved passages), and intra-memory conflict (inconsistent parametric beliefs), among which we focus on context-memory conflict. In this work, we focus on the context-memory conflict, which arises when a given information-bearing text chunk contradicts the model’s parametric beliefs.
Nuanced Behaviors under Conflict
Early studies reported that models tend to rely on their own knowledge when the prompt provides contradictory evidence (Longpre et al., 2021; Chen et al., 2022). Later work revealed a more nuanced picture. On synthetic datasets, Xie et al. (2023) showed that LLMs often update their answers when given strong and convincing evidence, whereas Jin et al. (2024a) observed a “Dunning–Kruger” effect in stronger LLMs, which display higher confidence in their incorrect parametric knowledge than in the external context. Further analysis also finds that models show availability bias (leaning on common-knowledge facts), majority bias (trusting the answer supported by more frequent evidence across documents), and confirmation bias (preferring evidence consistent with their prior knowledge), especially when the models are given misleading or irrelevant answers. Moving to realistic documents, Kortukov et al. (2024) found that models update their answers more reliably than synthetic evaluations suggest, yet still exhibit a parametric bias: if the model’s originally believed answer appeared anywhere in the context (even as a distractor), the model was more likely to stick to that incorrect answer.
Mitigation Strategies
Methods have also been proposed to alleviate context-memory knowledge conflict. Jin et al. (2024b) identified certain attention heads that specialize in “memory” while others specialize in “context”, and therefore propose a method that dynamically prunes or patches specific attention heads that cause conflicts. Efforts have also been made to develop novel decoding methods that enhance the use of contextual knowledge Jin et al. (2024a); Shi et al. (2024); Wang et al. (2025).
Our Focus
Most prior studies focus on contextual question answering, a setting that requires heavy reliance on the provided passages. Many other tasks, for example, grammar correction or claim verification, may need little context or, conversely, require careful integration of both parametric and contextual knowledge. This leaves the question of whether context-memory conflict poses the same impact on tasks with different knowledge demands, unanswered. To fill this gap, we keep the underlying knowledge constant while varying the task formulation, creating controlled datasets that induce different conflict levels for each target model. We introduce an analysis tool that automatically constructs model-specific test sets. Our findings indicate that both knowledge-memory conflict and blindly following the context could be particularly harmful to model-based evaluations and defending prompt injections.
@@ 3 Context-Memory Conflict Creation
Figure 2 illustrates an overview of the data construction pipeline. The process begins with identifying the pre-existing knowledge within a language model (Parametric Knowledge Collection). We use knowledge in question answering datasets that have two or more acceptable answers to one question, designed to study knowledge conflict Wan et al. (2024); Hou et al. (2024), as the knowledge source to find the stance that matches the model’s parametric belief, which will further be used to create task data. A piece of knowledge is considered part of the model’s internal belief only if the model consistently aligns with the perspective in a single answer across all prompt variations under greedy decoding, while rejecting conflicting alternatives. The variation of prompts is included in Appendix A.
With the model’s internal knowledge established, the framework generates contradictory statements based on a spectrum of conflict levels (§3.1, Evidence Creation). Leveraging these controlled contradictions, we build diagnostic datasets that consist of tasks requiring contextual knowledge, parametric knowledge, or a combination of both (§3.2, Task-Annotation). Since different models possess different parametric knowledge, the exact knowledge included in the diagnostic datasets differs by model. Each instance is then reviewed by an LLM to verify the correctness of its task type annotation (Validation).
@@ 3.1 Evidence Creation
The cognitive science literature suggests that humans address conflict between their knowledge and new information through cognitive judgment of the rationality of the concept Posner et al. (1982); Vosniadou and Brewer (1992). Xie et al. (2023) in turn suggests that LLMs could also update their answer if the provided context is convincing. We formalize it with the notion of plausibility to create a fine-grained taxonomy of conflict levels. Plausibility is defined as “at a minimum, the individual is willing to consider an alternative strategy because the recommendation is understood, coherent, and relatively simple and because the proposal is deemed a viable and logical alternative to solve the specific challenge at hand” Posner and Strike (1992). Plausibility can be used to measure how likely a human would accept new information when conflict exists. We quantify this notion by decomposing plausibility into two aspects: the content aligns with real-world or commonsense knowledge and does not violate basic logical principles. For example, suppose the model believes that grooves on the surface of Phobos, a moon of Mars, were caused by a boulder from an asteroid ejection. The conflicting statement that it was caused by gravitational pull from Mars is plausible because it conforms to common-sense knowledge. However, the idea that it was caused by a dance party is of low plausibility. With this in mind, we define three types of instances based on their alignment with the model’s internal knowledge (fig.˜1): No Contradiction (NC), High Plausibility Contradiction (HPC), Low Plausibility Contradiction (LPC).
@@
The evidences are created following fig.˜2. Starting with an original dataset $D_{\text{orig}}=\{(q_{i},\{a_{i1},a_{i2},...\},\{c_{i1},c_{i2},...\}),i\in[1,N]\}$, where $q_{i}$, $a_{i}$, $c_{i}$ corresponds to the question, answer, and context (supporting passage) of the $i$-th instance, $N$ is the size of dataset $D_{\text{orig}}$. The subscript $j$ after $i$ represents the $j$-th answer/context of the question $q_{i}$, as each question $q_{i}$ may have multiple acceptable answers. Since $D_{\text{orig}}$, coming from ConflictQA and WikiContradict, contains realistic and factually verified answers and contexts, we treat these existing answers as highly plausible. When an answer $a_{ij}$ from the original dataset contradicts the model-aligned answer $a_{ik}$ in an NC instance, we designate it as an HPC answer ($a_{i}^{\text{HPC}}=a_{ij}$), and its corresponding context as an HPC passage ($p_{i}^{\text{HPC}}=c_{ij}$). The contradicting answer $a_{ik}$ therefore becomes the NC example, namely, $a_{i}^{\text{NC}}=a_{ik}$ and $p_{i}^{\text{NC}}=c_{ik}$. To generate additional variants, we pass the passage $p_{i}^{\text{NC}}$ into an editor LLM, which is prompted to modify or rewrite it to achieve specified levels of plausibility and explanatory depth. Specifically, the editor model is instructed to rewrite the passage and degrade the plausibility while preserving contradiction to construct LPC passage $p_{i}^{\text{LPC}}$ and answer $a_{i}^{LPC}$. At the end of evidence creation, two LLMs were used to check (1) whether the passage-answer combination $(p_{i}^{\text{LPC}},a_{i}^{\text{LPC}})$ correctly answers the original question $q_{i}$; and (2) whether the generated context $p_{i}^{\text{LPC}}$ is truly low-plausibility through fact checking process.
@@ 3.2 Task Annotation
tasks do not require access to either contextual or parametric knowledge. We use extractive question answering as a KF task: the model is expected to extract a one-sentence answer directly from the context $p_{i}$ without engaging in reasoning, paraphrasing, or drawing upon prior knowledge. For example, the expected output in fig.˜1 should be “Grooves were formed during a massive dance party held by the witch among tiny alien creatures," which requires no additional change from the context. The list of acceptable extractions is obtained and verified by GPT-4o OpenAI (2024). In the evaluation setting, the output is treated as correct as long as the extracted sentence matches one of the acceptable extractions.
@@
One of the original datasets we use employs model-based evaluation to judge the correctness of free-text answers Hou et al. (2024). However, we observed that this evaluation method is susceptible to knowledge conflict, leading to inaccurate evaluations. We explore this issue further in §4.3. Therefore, we modify the non-extractive tasks to be multiple-choice questions. Each instance presents four answer options; the model must first generate an explanation, then select the most appropriate answer. To assess the performance of the target model, we report the accuracy for CK and PK tasks, F1 for KF, PCK, and RAG. To obtain high-quality texts, we use GPT-4o as the base model to create evidence and validate the diagnostic data. Then, we analyze the instruction-tuned version of Mistral-7B Jiang et al. (2023), OLMo2-7B, OLMo2-13B OLMo et al. (2024), Qwen2.5-7B, and Qwen2.5-14B Qwen et al. (2025), all of which are widely used open-weight models that represent diverse training paradigms. The resulting diagnostic data is composed of 2,893 instances for Mistral-7B, 177 instances for OLMo, and 6,217 instances for Qwen2.5-7B. Each instance includes three different evidence types (NC, HPC, LPC); thus, the resulting task data has three times the number of instances.